13.) A grating has 1000 lines/mm of width. How wide must the grating be in order to resolve the mode structure of a HeNe laser beam of wavelength = 633nm? Thefrequency difference between the modes is 450 MHz.
Answer #1

Since, the number of lines in 1mm is 1000. Thus, the grating element is -

d=\frac{0.1}{1000}cm=1\times 10^{-4}cm

and the number of lines is - N=1000

Therefore, the width of the grating is -

D=Nd=1000\times 10^{-4}=0.1 cm